OpenStudy (anonymous):

graph the equation by plotting 3 points -3y=-2x-6

OpenStudy (anonymous):

are helping me ?

OpenStudy (anonymous):

i mean are you helping me?

OpenStudy (anonymous):

take x=0, this gives -3y = 0 - 6 y = -6/-3 = 2 so (x,y) = (0,2) take x = -3, this gives -3y = 6 - 6 y = 0/-3 = 0 so (x,y) = (-3,0) take x = 3, this gives -3y = -6 -6 y = -12/-3 y = 4 so (x,y) = (3,4)

OpenStudy (anonymous):

plot the three points and join them with a line this will give the graph of the equation

OpenStudy (anonymous):

THANKKKKKKK YOU SOOOOOO MUCHHHHHHHH CAN YOU PLEASE HELP ME AGAIN =)))))))))!!!!!!!!!!!!

OpenStudy (anonymous):

go ahead and ask

OpenStudy (anonymous):

same thing as the last one

OpenStudy (anonymous):

-2x-3y=-7

OpenStudy (anonymous):

rearranging we get -3y = 2x -7 put x = -1 -3y = -2 -7 y = -9/-3 = 3 so (x,y) = (-1,3) put x = 2 -3y = 4 - 7 y = -3/-3 = 1 so (x,y) = (2,1) put x = 5, you get -3y = 10-7 y = 3/-3 = -1 so (x,y) = (5,-1) plot the three points to get your graph....

OpenStudy (anonymous):

pls click GOOD ANSWER button ( :})

OpenStudy (anonymous):

ok